OpenStudy (anonymous):

ln Sqrt e= 1/2 I was wondering how to put this in exponential form

OpenStudy (anonymous):

sqrt is same as ^(1/2)

OpenStudy (anonymous):

\[\LARGE \ln\sqrt e=\frac12\] \[\LARGE \sqrt{e}=\left(e\right)^{\frac12}\] lol...